The sum of the first 12 terms of an AP is 222 and the sum of the first 5 terms is 40. Write out the of the first four terms of the series

     


1
Expert's answer
2021-01-22T14:19:18-0500

"S_{12}=222, and , S_5=40"

"S_n=n\/2(a+(n-1)d)"

at "n=12"

"S_{12}=12\/2(a+(12-1)d)=6(a+11d)=222"

"a+11d=37 .........(i)"

at "n=5, S_5=5\/2(a+4d)=40"

"a+4d=(40\times2)/5"

"a+4d=16.........(ii)"

from (i)

"37-11d+4d=16"

"37-7d=16"

"7d=37-16=21"

"d=3"

from (i)

"a=37-11d=37-11\times3=4"

hence the first four terms are;

"a, a+d, a+2d, a+3d"

"4, 4+3, 4+2\times 3, 4+3\times 3"

"4, 7, 10, 13"